# RiNALMo: General-Purpose RNA Language Models Can Generalize Well on Structure Prediction Tasks

### v1

@@ Abstract
Ribonucleic acid (RNA) plays a variety of crucial roles in fundamental biological processes. Recently, RNA has become an interesting drug target, emphasizing the need to improve our understanding of its structures and functions. Over the years, sequencing technologies have produced an enormous amount of unlabeled RNA data, which hides important knowledge and potential. Motivated by the successes of protein language models, we introduce RiboNucleic Acid Language Model (RiNALMo) to help unveil the hidden code of RNA. RiNALMo is the largest RNA language model to date with $650$ million parameters pre-trained on $36$ million non-coding RNA sequences from several available databases. RiNALMo is able to extract hidden knowledge and capture the underlying structure information implicitly embedded within the RNA sequences. RiNALMo achieves state-of-the-art results on several downstream tasks. Notably, we show that its generalization capabilities can overcome the inability of other deep learning methods for secondary structure prediction to generalize on unseen RNA families. The code has been made publicly available on https://github.com/lbcb-sci/RiNALMo.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) trained on massive text corpora have been performing remarkably on various natural language understanding and generation tasks (Devlin et al., 2018; Radford et al., 2018; Raffel et al., 2020; Brown et al., 2020; Touvron et al., 2023; Chowdhery et al., 2023). In recent years, the exploration of language models has gone beyond the domain of natural language processing, reaching into the realms of biology and its data. A vast amount of sequenced protein data provided a ground for training protein language models, and since they have proven to be an extremely valuable asset in protein generative (Ferruz et al., 2022; Madani et al., 2023) and structure prediction tasks (Wu et al., 2022; Lin et al., 2023).
@@
Ribonucleic acids (RNAs) play crucial roles in fundamental biological processes, including transcription, cell signaling, chromatin remodeling, and genome imprinting. Like proteins, RNAs have recently become an attractive drug target, whose function and the way they interact with other biological molecules are closely related to their structure (Childs-Disney et al., 2022; Garner, 2023). However, much less attention has been given to applying language models to RNA-related problems, partly because there is no such amount of available data and corresponding structures, and partly because similar problems tend to be more difficult than for proteins.
@@
Currently, there are only two single-input-sequence RNA foundation models, RNA-FM (Chen et al., 2022) and Uni-RNA (Wang et al., 2023), that found applications and were tested on several structure and function prediction tasks. RNA-FM is a $100$ million parameters Transformer encoder based on the original implementation proposed by Vaswani et al. (2017) and trained exclusively on $23.7$ million non-coding RNAs (ncRNAs) from the RNAcentral database (RNAcentral Consortium, 2020). Wang et al. (2023) pre-trained an ensemble of language models ranging from $25$ to $400$ million parameters trained on a much larger dataset of $1$ billion sequences with the architecture analogous to the ESM protein language model (Rives et al., 2021). Unlike the Uni-RNA, RNA-FM is a publicly available RNA foundation model.
Figure 1: RiNALMo pre-training and applications. In the pre-training stage, RiNALMo is trained on unlabeled RNA sequences from several databases using masked language modeling (MLM). The final language model comprises $33$ Transformer blocks with the embedding dimension of $1280$. Once pre-trained, RiNALMo is separately fine-tuned for several structural and functional downstream tasks in which its expressive output embeddings, utilized by the prediction heads, significantly improve the performance.
@@
Motivated by the recent successes of protein language models and the latest architectural improvements in LLMs, we propose RiNALMo, a novel RNA language model. We pre-trained RiNALMo on a set of carefully curated $36$ million ncRNA sequences from the RNAcentral database augmented by several other RNA databases. RiNALMo is a $650$ million parameters BERT-style Transformer encoder (Devlin et al., 2018) advanced by modern architectural techniques such as rotary positional embedding (RoPE) (Su et al., 2024), SwiGLU activation function (Shazeer, 2020) and FlashAttention-2 (Dao, 2023). During pre-training, RiNALMo is able to extract hidden knowledge and capture the underlying structural information embedded within the sequences at the single-nucleotide level. Later, its output embeddings serve as a powerful sequence representation that, in comparison to other foundation models and state-of-the-art methods, improves the performance on various structural and functional RNA downstream tasks. In particular, RiNALMo shows remarkable generalization capability on secondary structure prediction of RNA families not encountered in the training dataset where other deep learning methods fail. RiNALMo pre-training procedure and downstream tasks are illustrated in Figure 1.
@@
  • We propose RiNALMo, a $650$ million parameters RNA language model, which is the largest RNA language model to date that can fully leverage the potential of a vast amount of public unannotated RNA sequences;
@@
  • We show that the generalization capability of RiNALMo can overcome the problem of other deep learning methods for secondary structure prediction to perform well on RNA families not seen in the training dataset;
@@
  • We conducted a set of extensive experiments on several RNA structural and functional downstream tasks whose results show that RiNALMo outperforms other RNA language models and deep learning methods on the majority of datasets.
The remainder of the paper is structured as follows. Section 2 presents the related work. In Section 2, we present RiNALMo and describe the dataset it was pre-trained on. In Section 4, we provide quantitative and qualitative assessments of RiNALMo evaluated on structural and several functional downstream tasks. In Section 5, we discuss the most important results and ideas of the paper. Section 6 concludes the paper.
2 Related Work
Protein language models.
Most efforts in applying the ideas originally developed for natural language processing have been focused on proteins after the success of AlphaFold (Jumper et al., 2021) in the prediction of protein structures. A vast amount of publicly available protein sequences has been seen as promising ground for studying generative and predictive models for protein-related biological problems. One of the first papers that applied self-supervised language modeling approaches to protein data was proposed by Rives et al. (2021). The authors’ LLM, ESM-1b, was pre-trained on $250$ million protein sequences and tested on several downstream tasks, including the secondary structure and tertiary contact prediction, where it achieved state-of-the-art results. Later on, several other protein LMs were proposed and tested on various downstream tasks (Heinzinger et al., 2019; Nambiar et al., 2020; Brandes et al., 2022; Elnaggar et al., 2022). Protein language models also play an important role in protein tertiary structure prediction. ESM-2 and OmegaPLM are examples of protein LMs that efficiently replace a multiple sequence alignment (MSA) step in ESMFold (Lin et al., 2023) and OmegaFold (Wu et al., 2022), which are deep learning-based structure prediction methods. Furthermore, several papers (Ferruz et al., 2022; Madani et al., 2023; Nijkamp et al., 2023; Chen et al., 2024) proposed language models that can generate protein sequences with specific biophysical properties. Recently, McWhite et al. (2023) showed that protein language models are a powerful new tool for generating MSAs, where the model’s output embeddings are leveraged to circumvent the need for many standard components of MSA algorithms.
RNA foundation models.
Chen et al. (2022) proposed RNA-FM, a BERT-style foundation model based on the original implementation of a Transformer encoder described in (Vaswani et al., 2017). RNA-FM is a $100$ million parameters model pre-trained on $23.7$ million unannotated ncRNAs from the RNAcentral database using the standard masked language modeling (MLM) task. It was used and tested on several structural and functional prediction tasks and found its application in an RNA 3D structure prediction method E2Efold-3D (Shen et al., 2022). A more recent RNA foundation model, Uni-RNA (Wang et al., 2023), was pre-trained on a dataset of $1$ billion RNA sequences for a spectrum of structural and functional downstream tasks. The model is a BERT-style Transformer encoder enhanced by several advanced techniques such as RoPE and fused layer norm. The authors pre-trained language models of different sizes and reported when the model parameters exceeded $400$M, the performance in downstream tasks reached a plateau with their architectures and datasets. Zhang et al. (2024) proposed an MSA-based RNA language model (RNA-MSM), a BERT-style language model that, instead of a single input sequence, utilizes a set of homologous sequences. RNA-MSM uses RNAcmap (Zhang et al., 2021) for an automatic homology search. Its embedding module consists of two position-embedding layers that encode rows and columns of the MSA separately. The authors showed that RNA structure information is preserved in the output attention maps and embeddings of RNA-MSM and that such an approach improves secondary structure prediction. However, obtaining the MSAs is a very time-consuming procedure – it takes RNAcmap on average $9$ hours to obtain an MSA for one RNA sequence of length $60$ (Zhang et al., 2024).
Language models for specific genomics-related downstream tasks.
Besides the foundation models, several authors proposed language models aimed at solving a few specific downstream tasks. Chen et al. (2023) proposed SpliceBERT, a BERT-style encoder pre-trained exclusively on more than $2$ million precursor messenger RNA (pre-mRNA) sequences from different vertebrates for studying RNA splicing. SpliceBERT outperforms DNABERT (Ji et al., 2021), a language model trained only on a human genome for prediction of promoters, splice sites and transcription factor binding sites, both on human and non-human splice-site prediction tasks. It demonstrates the superior performance and better generalization capability of language models pre-trained on multiple species. Later, the authors of DNABERT introduced DNABERT-2 (Zhou et al., 2023), which is a more advanced language model trained on a multi-species genome dataset. Yang et al. (2022) proposed single-cell BERT (scBERT), a BERT-style language model pre-trained on huge amounts of unlabeled single-cell RNA-seq data for cell type annotation. BigRNA (Celaj et al., 2023) is a language model consisting of an ensemble of $7$ models pre-trained on the genomes of $70$ individuals on a task to predict DNA-matched RNA-seq data. BigRNA accurately predicts tissue-specific RNA expression and the binding sites of proteins and microRNAs.
3 RNA Language Model
Figure 2: t-SNE visualization of sequence embeddings produced by RiNALMo for RNAs from the inter-family generalization evaluation dataset.
Our language model is an encoder-only Transformer that focuses on the task of understanding and unveiling the RNA code. At the heart of the model is the self-attention mechanism (Vaswani et al., 2017), which allows the capturing of important local and global contextual information. Before being passed to the Transformer, an RNA sequence is tokenized and turned into a $1280$ dimension vector using a learned input embedding model. RiNALMo consists of $33$ Transformer blocks, and each block comprises a multi-head attention and a feed-forward network (FFN). The position of the tokens is encoded using the RoPE (Su et al., 2024), which effectively encapsulates both the relative and the absolute positional information. Each multi-head attention has $20$ attention heads, which scales the model to $650$M parameters. To improve the pre-training efficiency of such a large model, we employed the IO-aware FlashAttention-2 (Dao, 2023), a fast and memory-efficient exact attention. In the FFN, we use two linear layers and the SwiGLU activation function (Shazeer, 2020) which combines the advantages of the Swish activation function and the gated linear unit (GLU). The Transformer modules are interconnected using residual connections. We use the layer normalization put inside the residual blocks to stabilize the training and have well-behaved gradients at initialization (Xiong et al., 2020).
We pre-trained RiNALMo using the MLM where we corrupted unlabeled RNA sequences and then tasked the model to reconstruct them. In this paper, each nucleotide is a single token. To corrupt the input sequence, we randomly selected $15\%$ of the tokens in the training sequence. Of these, $80\%$ are masked, i.e., replaced with the unique vocabulary token [MASK], $10\%$ are replaced with a randomly selected token from the vocabulary, and the remaining $10\%$ are left intact. We used ordinary cross-entropy training loss function. RiNALMo’s pre-training procedure and architecture are shown in Figure 1. More details on the pre-training and model’s hyperparameters can be found in Appendix A.
Besides using the RNAcentral database (RNAcentral Consortium, 2020), we collected sequences from Rfam (Kalvari et al., 2020), nt (Sayers et al., 2022) and Ensembl (Martin et al., 2022) as well, making it a pre-training dataset of $36$ million unique ncRNA sequences. To ensure sequence diversity in each training batch, we clustered the sequences with MMSeqs2 (Steinegger & Söding, 2017) into $17$ million clusters and then sampled each sequence in the batch from a different cluster.
Due to the computational limitations, during training, we limited the length of the input into the language model to $1024$ tokens. Every sequence begins with a [CLS] token, i.e., a classification token, and ends with an [EOS] token, i.e., an end-of-sequence token. If an input sequence is longer than $1022$ nucleotides, a random part of the sequence will be cropped out and fed into the model during each epoch, similar to what was done in (Rives et al., 2021).
@@ 3 RNA Language Model
Once pre-trained, RiNALMo’s output embeddings can serve as a powerful sequence representation that has embedded structural and evolutionary information. Such a representation can be used as an enriched input to structural and functional downstream tasks. We employed RiNALMo in a few prediction tasks to assess its performance and generalization capabilities. However, we anticipate it can be leveraged in many other tasks related to RNA structure and function.
@@
To qualitatively visualize pre-trained RiNALMo’s sequence representations, we employed t-SNE on the classification token embeddings for RNAs from a secondary structure prediction dataset (see Figure 2). In the output embedding space, it can be seen that the RNAs are clustered by families, and each family has a similar structure, with, in general, clean boundaries between clusters.
4 Downstream Tasks
We used pre-trained RiNALMo to solve one structure and two function prediction tasks. First, we show how RiNALMo can improve secondary structure prediction. In particular, RiNALMo’s generalization capability overcomes the inability of other deep learning methods for secondary structure prediction to generalize on unseen RNA families. Second, we show that RiNALMo can be successfully utilized in multi-species splice-site and mean ribosome loading prediction tasks where it achieves state-of-the-art results. In this section, we report the most important results on the downstream tasks, and more details and the ablation study can be found in Appendix B.
4.1 Secondary Structure Prediction
@@
When RNAs fold into complex structures, many of their bases pair up and form hydrogen bonds. These pairs are vital for the structure’s stability and function. Information about secondary structure alone can tell us a lot about the RNA and is often used as an input to the tertiary structure prediction tools.
@@
Most popular secondary structure prediction tools often rely on thermodynamic models, aiming to identify secondary structures that possess the lowest free energy (Reuter & Mathews, 2010). There are also popular probabilistic methods based on statistical learning procedures that act as an alternative to free energy minimization methods, such as CONTRAfold (Do et al., 2006). Several deep learning methods have been developed as well, and while they outperform the thermodynamic models on RNA families on which they were trained, they are usually unable to generalize well on new families. This is a severe limitation as it hinders practical usage of such tools (Szikszai et al., 2022).
We fine-tuned RiNALMo on a simple binary classification task with binary cross-entropy loss, where we tasked the model to classify each nucleotide pair as either paired or unpaired. A simple bottleneck residual neural network (ResNet) (He et al., 2016) is attached to the language model and is fed with nucleotide pair representations obtained by doing an outer concatenation of RiNALMo’s output with itself.
@@
First, we utilized datasets proposed in Singh et al. (2019) and compared our model to RNA-FM (Chen et al., 2022) and popular deep learning methods specialized for secondary structure prediction SPOT-RNA (Singh et al., 2019), UFold (Fu et al., 2021) and MXfold2 (Sato et al., 2021). Proposed training and evaluation datasets are denoted as TR0 and TS0 respectively and are non-redundant to each other on the sequence similarity level. All models were trained on the same training dataset (TR0) except SPOT-RNA which was additionally fine-tuned on a smaller dataset derived from PDB (Berman et al., 2000). Notice that sequences from the same RNA family can appear in both the training and evaluation datasets. As can be seen in Table 1, RiNALMo significantly outperforms other approaches.
Table 1: Performance of different models on the TS0 evaluation dataset.
Model Precision Recall F1
SPOT-RNA 0.70 0.62 0.64
UFold 0.63 0.71 0.66
MXfold2 0.67 0.57 0.60
RiNALMo 0.78 0.73 0.75
RNA-FM 0.71 0.66 0.68
@@
To test the generalization capabilities of our model, we utilized the dataset proposed in (Szikszai et al., 2022). The dataset of $3865$ RNAs from nine families was split nine times, and in each split, a different family was held out for evaluation while the other eight families were used for training and validation. RiNALMo’s ability to generalize across different RNA families was compared to RNA-FM (Chen et al., 2022), popular thermodynamics-based tool RNAstructure (Reuter & Mathews, 2010), probabilistic method CONTRAfold (Do et al., 2006), and two deep learning approaches specialized for secondary structure prediction UFold (Fu et al., 2021) and MXFold2 (Sato et al., 2021). The language models were fine-tuned and the other two deep learning models were separately trained on each of the previously described dataset splits and evaluated on a corresponding unseen RNA family. For predicting the secondary structures using CONTRAfold, we used EternaFold parameters (Wayment-Steele et al., 2022) trained on the EternaBench dataset, a set of more than $20,000$ RNAs.
F1 scores for different dataset splits are shown in Table 2. Fine-tuned RiNALMo demonstrates that it is more than capable of inter-family generalization as it outperforms RNAstructure in eight out of nine families by high margins, unlike other deep learning models. It is worth noting, however, that RiNALMo struggles to generalize on telomerase RNAs, but it achieves the highest F1 score on all other families. Visualization of RiNALMo’s sequence embeddings for all RNAs in the dataset is presented in Figure 2. One can notice that telomerase RNAs are clustered together, however, there is no clear boundary between them and SRP RNAs. We also noticed that telomerase RNAs are the longest in the dataset, on average around $25\%$ longer than the second-longest in the dataset. Interestingly, UFold performs best on telomerase RNA, while achieving much worse results on the other families. We are currently unable to conclude why RiNALMo fails on telomerase RNAs, but will take more focus on this problem in the future.
@@
Table 2: Inter-family generalization evaluation F1 scores for secondary structure prediction.
Test Family RNAstructure CONTRAfold RiNALMo RNA-FM MXfold2 UFold
5S rRNA 0.63 0.63 0.88 0.57 0.54 0.53
SRP RNA 0.63 0.55 0.70 0.25 0.50 0.26
tRNA 0.70 0.77 0.93 0.79 0.64 0.26
tmRNA 0.43 0.49 0.80 0.28 0.46 0.40
RNase P RNA 0.55 0.63 0.80 0.31 0.51 0.41
Group I intron 0.54 0.60 0.66 0.16 0.45 0.45
16S rRNA 0.57 0.58 0.74 0.14 0.55 0.41
Telomerase RNA 0.50 0.54 0.12 0.07 0.34 0.80
23S rRNA 0.73 0.71 0.85 0.19 0.64 0.45
Mean 0.59 0.61 0.72 0.31 0.51 0.44
When assessing the performance of secondary structure prediction, it is important to consider RNA structural dynamics, and that’s why it is helpful to view predictions that are close enough as correct. We employed the metric calculation approach proposed in Mathews (2019). To be more precise, for a nucleotide pairing $(i,j)$ where $i$ and $j$ represent nucleotide indices in the RNA sequence, $(i\pm 1,j)$ and $(i,j\pm 1)$ pairings are also considered correct predictions. The same approach has been utilized in both Table 1 and Table 2.
An example of secondary structure prediction from the inter-family generalization task can be seen in Figure 3. More details on the pipeline and examples of the secondary structure prediction can be found in Appendix B.1.
Figure 3: An example of secondary structure prediction. The shown secondary structure is a member of the 5S rRNA family and was used as a part of the evaluation dataset. RiNALMo is able to reconstruct the secondary structure accurately despite being fine-tuned with the dataset without any 5S rRNA. More examples of secondary structure predictions can be found in Appendix B.1 in Figure 5.
4.2 Multi-Species Splice-Site Prediction
@@
RNA splicing plays an important role in eukaryotic gene expression, involving the removal of introns from precursor messenger RNAs (pre-mRNAs) and the ligation of exons to form mature mRNAs. Precisely pinpointing splice sites—the donor and acceptor sites that mark the boundaries between exons and introns, and vice versa—is essential for predicting gene structure and location accurately.

### v2

@@ Abstract
While RNA has recently been recognized as an interesting small-molecule drug target, many challenges remain to be addressed before we take full advantage of it. This emphasizes the necessity to improve our understanding of its structures and functions. Over the years, sequencing technologies have produced an enormous amount of unlabeled RNA data, which hides a huge potential. Motivated by the successes of protein language models, we introduce RiboNucleic Acid Language Model (RiNALMo) to unveil the hidden code of RNA. RiNALMo is the largest RNA language model to date, with 650M parameters pre-trained on 36M non-coding RNA sequences from several databases. It can extract hidden knowledge and capture the underlying structure information implicitly embedded within the RNA sequences. RiNALMo achieves state-of-the-art results on several downstream tasks. Notably, we show that its generalization capabilities overcome the inability of other deep learning methods for secondary structure prediction to generalize on unseen RNA families.
@@ 1 Introduction
Large language models (LLMs) trained on massive text corpora have been performing remarkably on various natural language understanding and generation tasks [1, 2, 3, 4, 5, 6]. In recent years, the exploration of language models (LMs) has gone beyond the domain of natural language processing (NLP), reaching into the realms of biology and its data. A vast amount of sequenced protein data provided a ground for training protein LMs, and since they have proven to be an extremely valuable asset in protein generative [7, 8, 9] and structure prediction tasks [10, 11].
Most efforts in applying the ideas originally developed for NLP have been focused on proteins after the success of AlphaFold [12] in the prediction of protein structures. ESM-1b [13] was one of the first models that applied the self-supervised language modeling approaches to protein data. It was pre-trained on $250$M protein sequences and tested on several downstream tasks, including the secondary structure and tertiary contact prediction, where it achieved state-of-the-art results. Later on, several other protein LMs were proposed and tested on various downstream tasks [14, 15, 16, 17]. Protein LMs play an important role in protein tertiary structure prediction. ESM-2 [11] and OmegaPLM [10] are examples of protein LMs that efficiently replace a multiple sequence alignment (MSA) step in deep learning (DL) methods for structure prediction.
@@
RNAs play crucial roles in fundamental biological processes, including transcription, cell signaling, chromatin remodeling, and genome imprinting. Like proteins, RNAs have recently become an attractive drug target, whose function and interaction with other molecules are closely related to their structure [18, 19]. However, much less attention has been given to applying LMs to RNA-related problems, partly because there is no such amount of available data and corresponding structures, and partly because similar problems tend to be more difficult than for proteins.
@@
Currently, there are only two single-input-sequence RNA foundation models, RNA-FM [20] and Uni-RNA [21], that found applications in several structure and function prediction tasks. RNA-FM is a $100$M parameters Transformer encoder based on the original implementation by Vaswani et al. [22] and trained exclusively on $23.7$M non-coding RNAs (ncRNAs) from the RNAcentral database [23]. Wang et al. [21] pre-trained an ensemble of LMs ranging from $25$M to $400$M parameters trained on a much larger dataset of $1$B sequences with the architecture analogous to the ESM protein LM [13] enhanced by several advanced techniques such as RoPE and fused layer norm. The authors pre-trained language models of different sizes and reported when the model parameters exceeded $400$M, the performance in downstream tasks reached a plateau with their architectures and datasets. Both foundation models used the standard BERT-style masked language modeling (MLM) pre-training task [1]. Unlike the Uni-RNA, RNA-FM is publicly available.
Besides the RNA-FM and Uni-RNA foundation models, several authors proposed LMs to solve a few specific downstream tasks. RNA-MSM [24], an MSA-based BERT-style RNA LM, specialized in particular for secondary structure prediction, that instead of a single input sequence utilizes a set of homologous sequences. However, obtaining the MSAs is a very time-consuming procedure–it takes RNAcmap, a homology search tool used by RNA-MSM, on average nine hours to obtain an MSA for one RNA sequence of length $60$ [24]. Chen et al. [25] proposed SpliceBERT, a BERT-style encoder pre-trained exclusively on more than $2$M precursor messenger RNA (pre-mRNA) sequences from different vertebrates for studying RNA splicing. SpliceBERT outperforms DNABERT [26], an LM trained only on a human genome, both on human and non-human splice-site prediction tasks. It demonstrates better generalization capability of LMs pre-trained on multiple species. Yang et al. [27] proposed single-cell BERT (scBERT), a BERT-style LM pre-trained on huge amounts of unlabeled single-cell RNA-seq data for cell type annotation. BigRNA [28] is an LM pre-trained on the genomes of $70$ individuals on a task to predict DNA-matched RNA-seq data. BigRNA accurately predicts tissue-specific RNA expression and the binding sites of proteins and microRNAs.
@@
Motivated by the recent successes of protein LMs and the latest architectural improvements in LLMs, we propose RiNALMo, a novel RNA language model. We pre-trained RiNALMo on a set of carefully curated $36$M ncRNA sequences from the RNAcentral database augmented by several other RNA databases. RiNALMo is a $650$M parameters BERT-style Transformer encoder advanced by modern architectural techniques such as rotary positional embedding (RoPE) [29], SwiGLU activation function [30] and FlashAttention-2 [31]. During pre-training, RiNALMo can extract hidden knowledge and capture the underlying structural information embedded within the sequences at the single-nucleotide level. Later, its output embeddings serve as a powerful sequence representation that improves the performance on various structural and functional RNA downstream tasks compared to other foundation models and state-of-the-art methods. In particular, RiNALMo shows remarkable generalization capability on secondary structure prediction of RNA families not encountered in the training dataset where other DL methods fail.
@@
  • We propose RiNALMo, a $650$M parameters RNA LM, which is the largest RNA language model to date that can fully leverage the potential of a vast amount of public unannotated RNA sequences;
@@
  • We show that the generalization capability of RiNALMo can overcome the problem of other DL methods for secondary structure prediction to perform well on RNA families not seen in the training dataset;
@@
  • We conducted extensive experiments on several RNA structural and functional downstream tasks whose results show that RiNALMo outperforms other RNA LMs and DL methods on most datasets.
  • We release the pre-trained and fine-tuned RiNALMo weights and scripts for fine-tuning the model for the downstream tasks.
2 Results
[…]
General-purpose RNA language model A schematic diagram of RiNALMo and its pre-training procedure and downstream tasks are shown in Fig. 1.
Figure 1: RiNALMo pre-training and applications. In the pre-training stage, RiNALMo is trained on unlabeled RNA sequences from several databases using masked language modeling (MLM). To corrupt the input sequence, we randomly mask $15\%$ of the tokens in the training sequence. Before being passed to the Transformer, an RNA sequence is tokenized and turned into a $1280$ dimension vector using a learned input embedding module. The language model comprises $33$ Transformer blocks. Each Transformer block consists of a multi-head attention and a feed-forward network. Once pre-trained, RiNALMo can be separately fine-tuned for various structural and functional downstream tasks in which its expressive output embeddings, utilized by the prediction heads, significantly improve the performance. In this work, we fine-tuned RiNALMo for secondary structure, multi-species splice-site and mean ribosome loading (MRL) prediction tasks.
Our LM is a Transformer encoder focused on understanding and unveiling the RNA code. At the heart of the model is the self-attention mechanism [22], which captures important local and global contextual information. We pre-trained RiNALMo using the MLM where we tasked the model to reconstruct corrupted unlabeled RNA sequences. In this paper, each nucleotide is a single token. To corrupt the input sequence, we randomly mask $15\%$ of the tokens in the training sequence. To reconstruct the masked tokens, RiNALMo’s embeddings are utilized by the MLM prediction head whose outputs are used in the cross-entropy loss function. More technical details, pre-training details and ablation study are given in Methods and Supplementary materials.
@@ 2 Results
Once pre-trained, RiNALMo’s output embeddings can serve as a powerful sequence representation that has embedded structural and evolutionary information. First, its embeddings can be used for visualization and clustering analysis of RNA sequences. Second, such a representation can be used as an enriched input to structural and functional downstream tasks. We employed RiNALMo in a few tasks to assess its performance and generalization capabilities. Namely, we show how RiNALMo can improve and generalize well on secondary structure, multi-species splice-site and mean ribosome loading (MRL) prediction tasks. However, we anticipate it can be leveraged in many other tasks related to RNA structure and function. Particularly interesting would be the employment of RiNALMo in RNA tertiary structure prediction tasks where, motivated by the results from ESMFold [11] and OmegaFold [10], we believe RiNALMo’s embeddings can successfully replace the MSA.
@@
To visualize pre-trained RiNALMo’s sequence representations, we applied t-SNE on the classification token embeddings for RNAs from a secondary structure prediction dataset (see Fig. 2).
Figure 2: t-SNE visualization of sequence embeddings outputted by RiNALMo for RNAs from the inter-family generalization evaluation dataset.
In the embedding space, the RNAs are clustered by families with, in general, clean boundaries between clusters. This shows RiNALMo’s ability to cluster and distinguish different RNA families and confirms it can be used in various clustering analyses of RNA sequences. The t-SNE visualization is important from the RNA structure perspective as well since the RNAs from the same families fold similarly, meaning the structure information is also contained in the embedding space.
[…]
@@
Fine-tuning RiNALMo for intra-family secondary structure prediction When RNAs fold into complex structures, many of their bases pair up and form hydrogen bonds. These pairs are vital for the structure’s stability and function. These bonds can be represented by secondary structure, which can tell us a lot about RNA and which is often used as an input to the tertiary structure prediction tools. An example of a secondary structure can be seen in Fig. 3a.
Figure 3: Intra-family secondary structure prediction. a, RNAs fold into various shapes according to their function and while doing so, many of their nucleotides pair up using a hydrogen bond. These pairings are crucial for structure stability and form structural motifs such as hairpin loops and bulges. b, RiNALMo produces nucleotide embeddings for the given RNA sequence. Nucleotide pair embeddings are constructed by applying outer concatenation to RiNALMo’s outputs. Finally, pair representations are fed into the convolutional bottleneck residual neural network (ResNet) which produces base pairing probabilities that are then converted into the final secondary structure prediction. c, Precision, recall and F1 performance of different deep learning models on the TS0 evaluation dataset. d, Distribution of F1 scores for predictions of different models on the TS0 dataset. e, A target RNA from the TS0 evaluation dataset and its predictions from different deep learning models.
@@
Most popular secondary structure prediction tools often rely on thermodynamic models, aiming to identify secondary structures that possess the lowest free energy [32]. There are also popular probabilistic methods based on statistical learning procedures that act as an alternative to free energy minimization methods, such as CONTRAfold [33]. Several DL methods have been developed as well. They often outperform the thermodynamic models on RNA families on which they were trained, i.e., on in-distribution data.
@@
We fine-tuned RiNALMo on a simple binary classification task with binary cross-entropy loss, where we tasked the model to classify each nucleotide pair as either paired or unpaired. The pipeline for determining secondary structures is illustrated in Fig. 3b. We utilized a dataset proposed in [34] and compared our model to RNA-FM [20] and popular DL methods specialized for secondary structure prediction SPOT-RNA [34], UFold [35] and MXfold2 [36]. Proposed training and evaluation datasets are denoted as TR0 and TS0, respectively, and are non-redundant to each other on the sequence similarity level. However, notice that sequences from the same RNA family can appear in both the training and evaluation datasets. All models were trained on the same training dataset (TR0) except SPOT-RNA which was additionally fine-tuned on a smaller dataset derived from PDB [37]. As can be seen in Fig. 3c, RiNALMo outperforms other state-of-the-art DL approaches in terms of precision, recall and consequently F1 score. We provide F1 score distributions in Fig. 3d. A TS0 target example and the predictions from different DL methods are given in Fig. 3e.
[…]
RNA language models can generalize well on inter-family structure prediction tasks While DL methods for secondary structure prediction outperform thermodynamic models on in-distribution data, they are usually unable to generalize well on new RNA families. This is a severe limitation as it hinders practical usage of such tools [38].
@@
To test the generalization capabilities of RiNALMo, we utilized the dataset proposed in [38]. The dataset of $3865$ RNAs from nine families was split nine times, and in each split, a different family was held out for evaluation while the other eight families were used for training and validation. RiNALMo’s ability to generalize across different RNA families was compared to RNA-FM [20], popular thermodynamics-based tool RNAstructure [32], widely used probabilistic method CONTRAfold [33], and two DL approaches specialized for secondary structure prediction UFold [35] and MXFold2 [36]. The LMs were fine-tuned and the other two DL models were separately trained on each of the previously described dataset splits and evaluated on a corresponding unseen RNA family. For predicting the secondary structures using CONTRAfold, we used EternaFold parameters [39] trained on the EternaBench dataset, a set of more than $20,000$ RNAs.
Average F1 scores and F1 score distributions for different dataset splits are shown in Fig 4.
@@
Figure 4: Inter-family secondary structure prediction. a, Secondary structure prediction average F1 scores for the ArchiveII evaluation datasets. b, Distribution of F1 scores for different methods on the ArchiveII evaluation datasets.
Fine-tuned RiNALMo demonstrates that it is capable of inter-family generalization as it outperforms RNAstructure and CONTRAfold in eight out of nine families by high margins, unlike other DL models. To the best of our knowledge, this is the first paper to show that LMs can generalize well on inter-family secondary structure prediction, mitigating the limitations of other DL methods. We noted, however, that RiNALMo struggles to generalize on telomerase RNAs, but it achieves the highest F1 score on all other families. Visualization of RiNALMo’s sequence embeddings for all RNAs in the dataset is presented in Fig. 2. One can notice that telomerase RNAs are clustered together, however, there is no clear boundary between them and SRP RNAs. We also noticed that telomerase RNAs are the longest in the dataset, on average around $25\%$ longer than the second-longest in the dataset. Interestingly, UFold performs best on telomerase RNA, while achieving much worse results on the other families. We are currently unable to conclude why RiNALMo fails on telomerase RNAs, but will take more focus on this problem in the future. Technical details and more secondary structure prediction results and examples can be found in Methods, Supplementary materials and Extended Data Fig. 1.
[…]
@@
Fine-tuning RiNALMo for splice-site prediction RNA splicing plays an important role in eukaryotic gene expression, involving the removal of introns from pre-mRNAs and the ligation of exons to form mature mRNAs (see Fig. 5a). Precisely pinpointing splice sites—the donor and acceptor sites that mark the boundaries between exons and introns, and vice versa—is essential for accurately predicting gene structure and location.